OpenStudy (anonymous):

What is the height of a rectangular based pyramid with a volume of three hundred cubic meters and base dimensions of fifteen meters and ten meters? 2meter 6 meter 18 meter 54 meter

OpenStudy (welshfella):

volume of this pyramid = 300= (1/3) * (15*10) * h solve for h

OpenStudy (welshfella):

i'll start it off for you 300 = (1/3) * 150 * h 300 = 50 * h so what is h ?

OpenStudy (anonymous):

It's 6

OpenStudy (anonymous):

ThankYou

OpenStudy (welshfella):

yes 6 is correct